OpenStudy (anonymous):

an object is dropped from rest.if it falls a distance during S1 during the 1st second and an additional distance S2 in the next second.the ratio of S1/S2 is a-1 b-3 c-2 4-5

OpenStudy (anonymous):

i think it should be b........... in first second its velocity increases frm 0 to g...... then applyin eq of motion s= 0.t - gt^2 --- 2 which becomes g --- 2 then in second second (it funny writing scond second) its velocity incrs frm g to 2g again applying eq of motion now rembr int velocity is g for the second second.......the final for the first second s = gt + gt^2 ----- 2 as time diff is 1 s = g + g ---- 2 s= 3g -------- 2 compare dist traveld and u get ur ratio :)

OpenStudy (anonymous):

\[ r(t) = h_0 + v_0 + {g t^2 \over 2}\]for this example, h_0 and v_0 are both 0. \[{s_1 \over s_2} = {{r(1)} \over {r(2)-r(1)}} = {{{g} \over {2}} \over {{4g \over 2}-{g \over 2}}} = {1 \over 3}\]Therefore, none of the above is correct. Most likely it's a typo and you swapped numerator and denominator when copying the problem...